A 72.0 kg swimmer jumps into the old swimming hole from a tree limb that is 3.55 m above the water.

A) Use energy conservation to find his speed just as he hits the water if he just holds his nose and drops in.

-> 8.34 m/s

B) Use energy conservation to find his speed just as he hits the water if he bravely jumps straight up (but just beyond the board!) at 2.55 m/s

C) Use energy conservation to find his speed just as he hits the water if he manages to jump downward at 2.55 m/s
Answer #1

OPE = Akie mgan = mv2 va J2004 -1 8-34mis ] mgon = {my2 + Śm vy? 9.84 3.55 = * 2.852 +10v2 vo 7.94 mis 9.8 X 3.55 = { (-2.55²